OpenStudy (anonymous):

Whats 120 cos 15?/:

OpenStudy (anonymous):

12

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

@ganeshie8 Help?

ganeshie8 (ganeshie8):

il give a hint, see if it helps :)

ganeshie8 (ganeshie8):

15 = 45-30

OpenStudy (anonymous):

hmmm?

OpenStudy (anonymous):

its equal?

OpenStudy (anonymous):

@ganeshie8 you look so much like your mom in that picture

ganeshie8 (ganeshie8):

\(120 \cos (15)\)

ganeshie8 (ganeshie8):

\(120 \cos (45 - 30)\)

OpenStudy (anonymous):

69

ganeshie8 (ganeshie8):

use the cosine difference formula

ganeshie8 (ganeshie8):

\(\cos(A-B) = \cos A \cos B + \sin A \sin B\)

OpenStudy (anonymous):

Okay let me see

OpenStudy (anonymous):

is it -7.5?

OpenStudy (anonymous):

A biker rides on an upward-sloped road inclined at a 15.0° angle with the horizontal. If he moves at a constant speed of 2.0 meters/second for 60 seconds, what is the magnitude of his horizontal displacement? Thats my problem?

OpenStudy (anonymous):

im lost...

ganeshie8 (ganeshie8):

oh I see... you're right ! horizontal displacement = 120 cos 15

ganeshie8 (ganeshie8):

you dont wanto use calculator is it ?

OpenStudy (anonymous):

but none of my answers match/:

OpenStudy (anonymous):

wait what? sorry

ganeshie8 (ganeshie8):

list your options, we will see

OpenStudy (anonymous):

A. 3 meters B. 116 meters C. 60 meters D. 130 meters

ganeshie8 (ganeshie8):

im getting the answer as 116 meters.

ganeshie8 (ganeshie8):

you're not allowed to use calculator to solve this is it ?

ganeshie8 (ganeshie8):

if calculator is allowed, u can use wolfram directly : http://www.wolframalpha.com/input/?i=120cos%2815%29

OpenStudy (anonymous):

we can but they didnt give me one..

ganeshie8 (ganeshie8):

click the wolfram link. you can use it as calculator

OpenStudy (anonymous):

Thank you soo much :)

ganeshie8 (ganeshie8):

np :) if you wanto derive cos(15) you can do that as well using the formula i gave above.... good luck :)